OpenStudy (anonymous):

state the domain and range of the function f(x)-arc3sin(x/2)

OpenStudy (anonymous):

Could you write it in equation writer please?

OpenStudy (anonymous):

\[f(x) = 3\sin^{-1} (x/2)\]

OpenStudy (anonymous):

domain of arcsine is from [-1,1]

OpenStudy (anonymous):

in your case you should have \[-1\leq \frac{x}{2}\leq 1\] so \[-2\leq x \leq 2\] is your domain

OpenStudy (anonymous):

could you explain how you got the domain? the answers at the back said that but i dont quite get it

OpenStudy (anonymous):

range of arcsine is \[[-\frac{\pi}{2},\frac{\pi}{2}]\] so your range is \[[-\frac{3\pi}{2},\frac{3\pi}{2}]\]

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

if you want to take the arcsine of some number, that number has to be bigger than or equal to -1 and less than or equal to 1 this is clear yes?

OpenStudy (anonymous):

Arcsine is defined for: \[\forall x \in [-1;1], \arcsin(\sin(\theta))=\theta \] \[\rightarrow -1 \le \frac{x}{2} \le 1 \iff -2 \le x \le 2\] Domain is therefore \[D=[-2;2]\]

OpenStudy (anonymous):

but this function says first divide by 2. it is \[3\sin^{-1}(\frac{x}{2})\] sp of you divide by 2 first you are allowed numbers that when divided by 2 are between -1 and . therefore you numbers (x itself) has to be between -2 and 2

OpenStudy (anonymous):

some typos in what i wrote, but i hope it is clear

OpenStudy (anonymous):

thank you, both of you guys. i'm still trying to digest it, but ill get it eventually.

OpenStudy (anonymous):

? What do you mean it says it first? x/2 is the value of a sine, and the value of a sine HAS to be in [-1;1]... yet \[x \in [-2;2] \iff \frac{x}{2} \in [-1;1]\] Which is why the domain is [-2;2]

OpenStudy (anonymous):

someone1348, I actually understood that. thanks guys

OpenStudy (anonymous):

hope it is clear what i mean. how to i evaluate \[3\sin^{-1}(\frac{x}{2})\] for a given number replacement for x? first divide it by 2 then take the inverse sine then multiply by 3. so "first divide by 2" yes?

myininaya (myininaya):

\[-1<sinx<1\] the range of sinx is [-1,1] so the domain of sine inverse of x is [-1,1]

OpenStudy (anonymous):

can you also explain how to get the range?

OpenStudy (anonymous):

The range is the interval from the lowest value to the highest value. The lowest value arcsine can take is -pi/2, the highest is pi/2 It's the image interval by the function of the domain interval.

myininaya (myininaya):

ok sinx is not 1 to 1 so we want sine inverse of x to exist so we need to restrict the domain of sine so that it is 1 to 1 so that the inverse does exist the restricted domain of sinx is [-pi/2,pi/2]

myininaya (myininaya):

so the range of sine inverse is [-pi/2,pi/2]

OpenStudy (anonymous):

In other words, \[\arcsin([-1;1])=[-\frac{\pi}{2};\frac{\pi}{2}]\]

OpenStudy (anonymous):

i know how it works for just sinx, but what happens when you have a 3 in front of it

OpenStudy (anonymous):

like, for f(x)=inverse 3sin(x/2) in particular

OpenStudy (anonymous):

right! that is why you have to multiply by 3 to get your range. here is a nice picture (ignore the second one) and you will see what the range is http://www.wolframalpha.com/input/?i=y%3D3+arcsin%28x%2F2%29

OpenStudy (anonymous):

Well then it all gets multiplied by 3.. I mean, if the range of arcsin is \[[-\frac{\pi}{2};\frac{\pi}{2}]\] Then the range of 3 times arcsin is \[[-\frac{3 \pi}{2};\frac{3\pi}{2}]\]

OpenStudy (anonymous):

OHHHHHHH I get it, thank you guys, really.

OpenStudy (anonymous):

since the range of arcsine is \[[-\frac{\pi}{2},\frac{\pi}{2}]\] the range of 3 arcsine is \[[-\frac{3\pi}{2},\frac{3\pi}{2}]\]